### v1

# Assessing the Uncertainty and Robustness of Object Detection Models for Detecting Stickers on Laptops

Abstract
Refurbishing laptops extends their lives while contributing to reducing electronic waste, which promotes building a sustainable future. To this end, the Danish Technological Institute (DTI) focuses on the research and development of several applications, including laptop refurbishing. This has several steps, including cleaning, which involves identifying and removing stickers from laptop surfaces. DTI trained six sticker detection models (SDMs) based on open-source object detection models to identify such stickers precisely so these stickers can be removed automatically. However, given the diversity in types of stickers (e.g., shapes, colors, locations), identification of the stickers is highly uncertain, thereby requiring explicit quantification of uncertainty associated with the identified stickers. Such uncertainty quantification can help reduce risks in removing stickers, which, for example, could otherwise result in damaging laptop surfaces. For uncertainty quantification, we adopted the Monte Carlo Dropout method to evaluate the six SDMs from DTI using three datasets: the original image dataset from DTI and two datasets generated with vision language models, i.e., DALL-E-3 and Stable Diffusion-3. In addition, we presented novel robustness metrics concerning detection accuracy and uncertainty to assess the robustness of the SDMs based on adversarial datasets generated from the three datasets using a dense adversary method. Our evaluation results show that different SDMs perform differently regarding different metrics. Based on the results, we provide SDM selection guidelines and lessons learned from various perspectives.
Keywords Uncertainty Quantification  $\cdot$ Robustness Evaluation  $\cdot$ Object Detection  $\cdot$ Deep Neural Network  $\cdot$ Empirical Evaluation
1 Introduction
The European Union’s Circular Economy Action Plan (CEAP) highlights the need for sustainable operations to promote circular economy processes and encourage sustainable consumption [1]. One essential activity is refurbishing electronic devices (e.g., laptops) to extend their lives, reduce electronic waste, and provide affordable options for consumers. A critical and time-consuming step in refurbishment is removing stickers from the laptop by first identifying stickers and their locations on the laptops. Manual cleaning is time-consuming and faces challenges in finding enough workers, and current automation solutions are not built for this amount of variation, limiting sustainability and scalability. Thus, novel solutions to transition from manual labor to automated processes are paramount. Robotics offers a promising solution to simplify and scale up this process, increasing efficiency and reducing labor costs.
The Danish Technological Institute (DTI) develops, applies, and transfers technology to industry and society. One leading area that DTI focuses on is laptop refurbishing automatically. Targeting the cleaning process in laptop refurbishment, DTI built several deep neural networks (DNN)–based sticker detection models (SDMs) for automatic sticker detection, which is the basis for successful automatic sticker removal. The SDMs are built on open-source object detection DNNs and trained using a sticker detection dataset specially designed by DTI. DNN has revolutionized various domains, such as computer vision [2] and autonomous systems [3], by providing powerful tools for predictive modeling and decision-making. However, due to inappropriate model architecture or insufficient training data, the design and training process may introduce uncertainties into the SDMs, making them vulnerable under certain conditions, such as adversarial attacks, noisy data, or unforeseen input distributions [4]. This vulnerability highlights the need to quantify uncertainty and evaluate the robustness of the SDMs, which is crucial for trustworthy sticker removal, as incorrect detection may damage the laptop surface or incomplete sticker removal.
Various approaches have been proposed to quantify DNN uncertainty. For example, Bayesian Neural Network (BNN) [5] provides a probabilistic framework for uncertainty quantification (UQ) through Bayesian inference, but it requires computationally complex posterior calculations. Monte Carlo Dropout (MC-Dropout) [6] is proposed as a Bayesian approximation based on dropout, offering a practical way to quantify prediction uncertainty by measuring the variability among multiple predictions. Regarding the robustness evaluation of DNNs, a typical technique is to evaluate the robustness under various adversarial attacks, referred to as adversarial robustness [7]. For instance, the Projected Gradient Descent (PGD) attack [8] assesses the adversarial robustness of DNNs from the robust optimization perspective.
In this paper, we conduct a comprehensive empirical evaluation to assess the SDMs that DTI uses in the laptop refurbishing process regarding detection accuracy, prediction uncertainty, and adversarial robustness. Specifically, we adopt MC-Dropout as the UQ method to capture the uncertainty in model predictions. We run the model to perform multiple predictions and based on these predictions, we calculate two types of UQ metrics: label classification UQ metrics and bounding box regression UQ metrics. We employ Dense Adversary Generation (DAG) [9] as the adversarial attack and define robustness score (RS) to measure the robustness of the SDMs. RS measures robustness from two perspectives: robustness concerning predictive precision and robustness concerning prediction uncertainty. Regarding benchmark datasets, we construct benchmark datasets from three data sources: datasets provided by our partner DTI, datasets synthesized by prompting two vision language models (VLMs), i.e., DALL-E-3 [10] and Stable Diffusion-3 [11], and datasets created using the adversarial attack, i.e., DAG. Our evaluation results show that different SDMs achieve different performance regarding different evaluation metrics. Specifically, regarding sticker detection accuracy, Faster R-CNN_v2 is recommended as the best SDM, while RetinaNet_v2 achieved the overall best performance regarding prediction uncertainty. Regarding adversarial robustness, Faster R-CNN_v2 and RetinaNet_v2 are recommended as the best SDMs. We also provide guidelines for SDM selection and lessons learned.
In summary, our contributions are: 1) we conduct an empirical study to evaluate the detection accuracy, uncertainty, and robustness of six SDMs in detecting stickers on the laptop; 2) we employ MC-Dropout as the UQ method and calculate UQ metrics to measure uncertainty in classification and regression; 3) we present a novel robustness metric, i.e., RS, to measure the robustness by considering model uncertainty; 4) we provide guidelines for SDM selection from different perspectives.
2 Industrial Context & Background
2.1 Industrial Context
DTI is a leading Danish institute focused on developing advanced technology and transferring the technology to industry. A key area that DTI focuses on is robotics. Within robotics, DTI is leading an initiative for robotic applications for sustainable operations deemed critical in the context of CEAP. Adopting circular economy principles from the CEAP could substantially boost the EU’s GDP by 0.5% by 2030, creating approximately 700,000 new jobs [1]. Electrical and electronic equipment, one of the fastest-growing waste streams in the EU, currently sees less than 40% recycled, resulting in significant value loss [1].
This scenario underscores the need for refurbishing electronics such as laptops with robots, which is the context of this paper. A typical refurbishment process consists of disassembly, cleaning, inspection, restoration, reassembly, and testing steps. Collaborative robots performing these steps often require human involvement for challenging tasks (e.g., non-trivial disassembly tasks). In particular, an essential and time-consuming step in the refurbishment process is cleaning the stickers on the laptop, which can be further divided into two tasks: sticker detection and removal. Accurate detection of stickers with high confidence is crucial, as incorrect detection may result in damage to the laptop surface or incomplete sticker removal. Although the sticker detection task is straightforward for humans, the unlimited number of stickers and ever-changing types of laptops pose a significant challenge for automating and scaling this process. This issue of handling extreme variability is common across many loops of the circular economy, including recycling, refurbishing, repairing, and remanufacturing. Therefore, having a method to manage this variability in a scalable and automated way is crucial for becoming more circular.
Robotics and AI systems can automate the sticker detection process. To this end, DTI leveraged computer vision technology and developed DNN-based SDMs to identify stickers on laptops, turning the sticker detection task into an automated process. During the DNN designing and training process, uncertainty might arise due to inappropriately designed model architectures or insufficient training data. For example, some laptop brands may not be included in the training data, so the SDMs may not detect stickers correctly or behave highly uncertainly (e.g., detecting the same object multiple times but giving different detection results). Therefore, we must quantify and manage the uncertainty in the refurbishment process of electronic products. Specifically, quantifying uncertainty in SDMs is crucial from several perspectives. First, it ensures safe and reliable decisions of robots, especially when robots work alongside humans. Besides, managing detection uncertainty can significantly improve quality control, reduce errors, and increase accuracy. In addition, adaptability can be enhanced by helping the model adapt to different sticker types and conditions, effectively guiding human intervention when needed. Finally, workflow efficiency can be optimized by prioritizing cases that require human attention, thereby reducing downtime and streamlining workflows.
Figure 1: Overview of the Uncertainty Quantification Process for Sticker Detection Models.
2.2 Bayesian Uncertainty Quantification in Deep Learning
Bayesian Neural Networks (BNNs) provide a probabilistic approach to neural network modeling that allows UQ through Bayesian inference [5]. In BNNs, model weights are treated as random variables with a prior distribution reflecting the initial belief about the weights before observing any data. A common choice is a Gaussian distribution: $p(\mathbf{w})=\mathcal{N}(\mathbf{w}\mid\mathbf{0},\mathbf{I})$. Given a dataset $\mathbf{X}=\{x_{1},x_{2},...,x_{n}\}$, $\mathbf{Y}=\{y_{1},y_{2},...,y_{n}\}$, Bayesian inference computes the posterior over the weights as: $p(\mathbf{w}\mid\mathbf{X},\mathbf{Y})$. Instead of providing a single output prediction, BNNs produce a distribution of possible outputs that captures the uncertainty about the prediction. The posterior combines the prior and the likelihood of the observed data, which is often difficult to compute exactly because the marginal probability $p(\mathbf{Y}\mid\mathbf{X})$ required to evaluate the posterior cannot be evaluated analytically [12]. Hence, various Bayesian approximation methods have been proposed to perform Bayesian inference.
Gal and Ghahramani [6] proposed MC-dropout as a Bayesian approximation method, which shows that introducing dropout in DNN can be interpreted as the approximation of a probabilistic Bayesian model in deep Gaussian processes. Dropout is commonly applied in deep learning models to avoid over-fitting by randomly dropping units along with their connections from the neural network [13]. To this end, the predictive distribution is approximated by sampling multiple predictions:
$p(y^{*}\mid x^{*},\mathbf{X},\mathbf{Y})\approx\frac{1}{T}p(y^{*}\mid x^{*},\mathbf{w}^{{}^{\prime}}),$ (1)
where $x^{*}$ is the input sample, $y^{*}$ is the corresponding predicted output, $T$ is the number of predictions, and $w^{*}$ is the model weights after dropout. The $T$ DNN predictions are different because neurons are randomly dropped in each prediction, based on which we quantify the uncertainty of the overall predictions.
3 Methodology
Figure 1 presents an overview of the UQ process for the SDMs. As the figure shows, the UQ process consists of dataset creation, MC-dropout-based prediction, and UQ metrics calculation. The dataset creation constructs benchmark datasets from three data sources: datasets from DTI, datasets synthesized by prompting two VLMs, and datasets created using an adversarial attack technique. For each image in the benchmark datasets, the SDM makes T predictions based on the MC-Dropout method to capture the prediction uncertainty. We then employ the density-based algorithm, HDBSCAN, to cluster the detected stickers in T predictions and calculate the UQ metrics based on the detected stickers. We consider two types of UQ metrics: label classification UQ metrics and bounding box regression UQ metrics. We present the MC-Dropout method in Section 3.1 and the UQ metrics in Section 3.2. Sections 3.3 and 3.4 introduce VLM-based and adversarial attack-based dataset creation.
3.1 Monte Carlo Dropout-Based Uncertainty Quantification
Bayesian probabilistic theory [14] has been shown as a universal theoretical framework for uncertainty reasoning, and as discussed in Section 2.2, MC-dropout is a Bayesian approximation for quantifying uncertainty. Hence, we adopt MC-Dropout as the UQ method and inject inference-time activated dropout layers into the pre-trained SDMs. By doing so, we build uncertainty-aware variants of the original models. Specifically, for a SDM $f$, we first build its variant as $f^{p}$ by applying dropout layers with dropout rate $p$. Next, we let $f^{p}$ make $T$ predictions for a given image input $X$ to obtain a sample of outputs $O=\{f^{p}_{t}\}_{t=1}^{T}$. The outputs $O$ represent a sample from the model’s predictive distribution, and MC-Dropout quantifies the uncertainty of $f_{p}$’s outputs by extracting information regarding the variability of the $T$ predictions. We set $T$ to 20 following the guidelines by Gal and Ghahramani [6].
3.2 Uncertainty Quantification Metrics
3.2.1 Basic Concepts
As discussed in Section 3.1, we utilize MC-Dropout to quantify the uncertainty over $T$ prediction outputs given an input image $X$: $O=\{f^{p}_{t}\}_{t=1}^{T}$. A prediction $f^{p}_{t}$ is a vector containing $N$ different detected objects, and a detected object $f^{p}_{t}(i),(1\leq i\leq N)$ is characterized by its label classification and bounding box regression. The label classification assigns the softmax score $sm_{f^{p}_{t}(i)}$ for the candidate objects of three classes: stickers, logos, and backgrounds. The softmax score is a vector of probabilities, and we take the position with the maximum probability value as the predicted label, i.e., $argmax()$. The bounding box regression returns four different real values $box_{f^{p}_{t}(i)}=\{x_{1},x_{2},y_{1},y_{2}\}$ to locate the detected object. Specifically, a bounding box is a rectangle whose start and end points are denoted as $(x_{1},y_{1})$ and $(x_{2},y_{2})$. In a deterministic model without uncertainty, all predictions should be the same, for example, a detected object $f^{p}_{t}(i)$ in prediction $f^{p}_{t}$ should also be detected in other predictions and have the same softmax score and bounding box regression values as those in other predictions. However, recall that MC-Dropout quantifies the uncertainty by utilizing $T$ model predictions with randomly dropped units. Therefore, each prediction is not necessarily the same. The model based on MC-Dropout can produce predictions with different locations for each object in a single image. To this end, a clustering method is needed to cluster objects in the $T$ predictions. Details will be introduced in the following section.
3.2.2 HDBSCAN-based Object Clustering
We adopt the hierarchical density-based clustering algorithm, HDBSCAN [15], to cluster objects based on their predicted bounding boxes. The algorithm identifies clusters by calculating core distances for each point and adjusting these distances based on local density. It then builds a minimum spanning tree to form a hierarchy of clusters, which is condensed to highlight meaningful clusters. This hierarchical structure enables the recognition of clusters of various shapes and sizes. Two key parameters need to be adjusted: minimum samples ($minSamples$), which sets the minimum number of samples in the neighborhood for a point to be considered a core point, and minimum cluster size ($minPts$), which determines the minimum number of points to be considered as a cluster. Based on our preliminary study, we set both $minSamples$ and $minPts$ to 3. As a result, each cluster represents a detected object in the $T$ predictions whose label classification softmax score and the bounding box regression are denoted as $\{sm_{k}\}_{k=1}^{W}$ and $\{box_{k}\}_{k=1}^{W}$. Note that $W$ is the number of times the object is detected, and it is in the range $[1,T]$, as some objects are not always detected in all $T$ predictions. We then calculate UQ metrics for each detected object in the $T$ predictions.
3.2.3 Uncertainty Metrics for Label Classification
We employ three common metrics to measure uncertainty in classification tasks [16], including Variation Ratio (VR[17], Shannon Entropy (SE[18], and Mutual Information (MI[18]. VR is a measure of dispersion. It is calculated by determining the proportion of cases that are not in the mode (the most frequent category) relative to the total number of cases. For the detected object in each cluster, we first extract a set of labels as the classes with the highest softmax scores:
$\mathbf{y}=\{\operatorname*{argmax}(sm_{k})\}_{k=1}^{W},$ (2)
we then find the mode of the distribution as:
$c^{*}=\operatorname*{argmax}_{c=1,...,C}\sum_{k}\mathds{1}[\mathbf{y}^{k}=c],$ (3)
where $\mathds{1}$ is the indicator function. Then the number of times $c^{*}$ was sampled is $\sum_{k}\mathds{1}[\mathbf{y}^{k}=c^{*}]$. Finally, VR is calculated as:
$\textit{VR}=1-\frac{\sum_{k}\mathds{1}[\mathbf{y}^{k}=c^{*}]}{W}.$ (4)
VR ranges from 0 to $2/3$. It attains the maximum value of $2/3$ when all three classes are sampled equally and the minimum value of 0 when only a single class is sampled.
SE captures the average amount of information contained in the predictive distribution. For a detected object, we calculate SE by considering softmax scores over $T$ predictions, i.e., $\{sm_{k}\}_{k=1}^{W}$ as:
$SE=-\sum_{c=1}^{nc}(\frac{1}{W}\sum_{k=1}^{W}sm_{k}(c))\times\log(\frac{1}{W}\sum_{k=1}^{W}sm_{k}(c)),$ (5)
where $nc$ denotes the number of classes, i.e., 3. SE values range from 0 to $\log(3)$, reaching a maximum of $\log(3)$ when all classes are predicted to have the same softmax score, i.e., $1/3$, indicating the most uncertain case. Its minimum value is 0 when the probability of one class is 1 and all other probabilities are 0, indicating no uncertainty in the classification.
MI quantifies the information difference between the predicted and posterior of the model parameters, providing a different uncertainty measure for classification tasks. For the detected object in each cluster, MI can be approximated as [16]:
$\begin{split}\textit{MI}=-\sum_{c=1}^{nc}(\frac{1}{W}\sum_{k=1}^{W}sm_{k}(c))\times\log(\frac{1}{W}\sum_{k=1}^{W}sm_{k}(c))\\ +\frac{1}{W}\sum_{k=1}^{W}\sum_{c=1}^{nc}sm_{k}(c)\times\log sm_{k}(c),\end{split}$ (6)
where $nc$ denotes the number of classes, i.e., 3. MI measures the model’s confidence in its outputs. It ranges from 0 to 1, and the larger the MI value, the higher the uncertainty.
Figure 2: An example for sticker detection. The left column shows the laptop image with two stickers on it. The right column shows the sticker detection outputs and what the convex hull looks like in the left image.
3.2.4 Uncertainty Metrics for Bounding Box Regression
For each detected object, we estimate the uncertainty for bounding box regression using Total Variance (TV[19] and Predictive Surface (PS[20]. TV captures the uncertainty in the bounding box regression by calculating the trace of the covariance matrix of $\{box_{k}\}_{k=1}^{W}$, which sums the variances of each variable in the bounding box. Specifically, a bounding box has four variables used to locate an object: $box=\{x_{1},x_{2},y_{1},y_{2}\}$, for variable $v\in box$, we calculate its variance as:
$\delta_{v}^{2}=\frac{1}{W-1}\sum_{k=1}^{W}(v_{k}-\mu_{v})^{2},$ (7)
where $\delta^{2}$ is the variance of $v$, $\mu$ is the mean value of $v$, and $W$ is the number of times the object is detected. We then calculate TV by summing the variances for all variables:
$\textit{TV}=\sum_{v\in box}\delta_{v}^{2}.$ (8)
TV ranges in $[0,+\infty)$, and a larger TV value indicates a higher uncertainty.
PS was originally proposed to quantify prediction uncertainty in object detection models for autonomous driving [20]. It measures uncertainty by considering the convex hull of each corner point of the predicted bounding box. A convex hull is the smallest convex shape enclosing all points and PS estimates uncertainty by calculating the area covered by the convex hull. To be concrete, four corner points define the bounding box to locate an object: $(x_{1},y_{1})$, $(x_{2},y_{2})$, $(x_{2},y_{1})$, and $(x_{1},y_{2})$. For the object in each cluster where the bounding box is detected $W$ times, i.e., $\{box_{k}\}_{k=1}^{W}$, we first obtain the cluster of each corner point as:
$\begin{split}CP=\{(x_{1},y_{1})_{k},(x_{2},y_{2})_{k},(x_{2},y_{1})_{k},(x_{1},y_{2})_{k}\}_{k=1}^{W},\end{split}$ (9)
we then identify the convex hulls for its four corner points, respectively, as:
$convexh=ConvexHull(cp),cp\in CP.$ (10)
Finally, we calculate PS by averaging the area covered by the convex hulls of the four corner points to approximate prediction uncertainty:
$\textit{PS}=\frac{1}{|CP|}\sum_{cp\in CP}Area(convexh_{cp}).$ (11)
PS ranges in $[0,+\infty)$, and a higher PS value indicates higher variability of the predicted corner points, meaning higher uncertainty in the bounding box prediction.
Recall that for an input image $X$, multiple objects will be detected, so after performing the HDBSCAN clustering algorithm, multiple clusters will be obtained, each cluster representing a detected object. To get the uncertainty of the input image $X$, for each UQ metric, we take the average of all metric values of all detected objects in $X$. Figure 2 shows an example of using the SDM to detect stickers on a laptop and how to calculate UQ metrics based on the model outputs. The SDM makes T predictions, and each prediction returns a label based on the softmax score and a bounding box containing four corner points. The convex hull of each corner point is then calculated. The UQ metrics for label classification are then calculated based on the softmax scores in T predictions. Regarding UQ metrics for bounding box regression, TV is calculated based on the variance of the values of each corner point, and PS is calculated based on the area covered by convex hulls.
3.3 VLM-based Image Generation
VLMs are generic pre-trained multimodal models that learn from large-scale image-text pairs sourced from the Internet, enabling them to directly address downstream visual tasks without task-specific fine-tuning [21]. Therefore, we apply the zero-shot generalization ability of VLMs to generate a synthetic image dataset to study whether SDMs can identify stickers on laptops that they have not encountered before. Following the representative categories of VLMs, i.e., autoregressive and diffusion models [22], we select two state-of-the-art VLMs: DALL-E-3 [10] and Stable Diffusion-3 [11]. DALL-E-3 is the latest VLM developed by OpenAI, which leverages advanced transformer architectures to generate images from textual descriptions. Stable Diffusion-3, developed by StablilityAI, is a state-of-the-art text-to-image generation VLM that uses diffusion processes to create high-quality images from textual descriptions.
To ensure consistency and equity in using VLMs, we design the same prompt template to prompt the selected VLMs. Table 1 lists the main prompt phrases and relevant examples, which involve common laptop models, the number, size, style, and location of generated stickers, and other detailed image content information. The prompt template consists of these phrases, e.g., "a laptop, three stickers on the lid, one small music sticker, one small animal sticker, one small superhero sticker, stickers close to each other, top-down view, screen turn off, on a flat surface, a solid dark background". Note that to enhance the generality of the synthesized dataset, the prompt template omits information about the quality of the generated images, e.g., sharp focus or highly detailed, thereby generating images with varying levels of clarity that can be used to test and evaluate the generalization capability of SDMs.
3.4 Adversarial Image Generation
DNNs are vulnerable to adversarial attacks, which are imperceptible to humans, but easily misclassified by DNNs [8]. To study the robustness of SDMs against adversarial attacks and how these attacks affect the prediction uncertainty of SDMs, we construct an adversarial image dataset by perturbing images in an original dataset. Specifically, given an image $X$ containing $N$ targets $\mathcal{T}=\{t_{1},t_{2},...,t_{N}\}$, each target is labeled by a ground-truth $l_{n}\in\{sticker,logo,background\}$, an adversarial attacker aims to inject perturbations on $X$ to obtain a perturbed image $X^{\prime}$ that satisfies:
$\forall 1\leq n\leq N\wedge t_{n}\in\mathcal{T}:f(X^{\prime},t_{n})=l_{n}^{\prime}\neq f(X,t_{n})=l_{n},$ (12)
where $f$ denotes the SDM. As a result, the generated adversarial image $X^{\prime}$ makes all targets incorrectly predicted. We adopt an adversarial attack approach called Dense Adversary Generation (DAG) [9] to construct the adversarial image dataset. DAG aims to generate adversarial examples for semantic segmentation and object detection tasks where multiple objects must be recognized. It requires an original dataset to perturb the images in it. We use the dataset to evaluate the pre-trained SDM as the original dataset and then add adversarial perturbations to the images in the original dataset to obtain an adversarial image dataset.
Table 1: Prompt Phrase with Example
Phrase Example
laptop model "a laptop", "a MacBook laptop", "a Dell laptop", "an HP laptop", …
sticker number "one", "two", "three", …
sticker size "small", "medium", "big"
sticker style "Apple logo", "Dell logo", …, "cartoon character", "animal", …
sticker location "close to the center of the laptop", "stickers close to each other", …
other "top-down view", "screen turn off", "a solid dark background", …
4 Experiment Design
4.1 Sticker Detection Models
We obtain six pre-trained SDMs from DTI, which represent state-of-the-art DNN-based object detection models. These models are designed by considering multiple model architectures which are Faster R-CNN [23], Faster R-CNN_v2 [24], RetinaNet [25], RetinaNet_v2 [26], SSD300 [27], and SSDLite [28]. DTI trained these models using a specially designed sticker detection dataset containing thousands of labeled laptop images with stickers in various poses, sizes, and lighting conditions. All six SDMs are trained based on the open-source implementation in PyTorch [29].
We modify the six pre-trained models to be compatible with the MC-Dropout method by injecting prediction-time activated dropout layers. Specifically, for Faster R-CNN, Faster R-CNN_v2, RetinaNet, and RetinaNet_v2, we add dropout layers to the convolutional layers of the Feature Pyramid Network, which is a common component of these four models and built on the backbones of these four models. Regarding SSD300 and SSDLite, we add dropout layers to their detection heads. In our experiments, we choose 9 dropout ratios from 0.1 to 0.5 with an interval of 0.05 to study the effect of dropout rates on the performance of SDMs.
4.2 Benckmark Datasets
We obtain a dataset containing 150 images from DTI, i.e., origImg, and synthesize two datasets by prompting two VLMs, DALL-E-3 and Stable Diffusion-3, i.e., dalleImg and stableImg, each containing 150 images. By applying the adversarial attack method (DAG) to origImg, dalleImg, and stableImg, we generate three adversarial datasets: advOrigImg, advDalleImg, and advStableImg. Specifically, for each image $X$ in each dataset, we execute DAG 10 times and obtain 10 adversarial images for $X$, and therefore, each adversarial dataset contains 1500 images. These three adversarial datasets are used to assess the robustness of the SDMs.
4.3 Evaluation Metrics
Mean Average Precision (mAP) is a commonly used metric for evaluating the accuracy of object detectors [30] and calculated as:
$\textit{mAP}=\frac{1}{nc}\sum_{c=1}^{nc}AP_{c},$ (13)
[…]
UQ Metrics (UQMs) are the UQ metrics defined in Section 3.2, which includes VR, SE, MI, TV, and PS.
Robustness Score (RS) measures the robustness of each SDM in performing the sticker detection task. We calculate RS regarding mAP and UQM, respectively, i.e., $RS_{mAP}$ and $RS_{uq}$ by considering the model’s performance on dataset $\mathcal{D}$ and $\mathcal{D}$’s adversarial version $\mathcal{D}_{adv}$. Specifically, for image $X\in\mathcal{D}$, there are 10 adversarial images $\{X_{adv}^{1},X_{adv}^{2},...,X_{adv}^{10}\}$ from $\mathcal{D}_{adv}$, and we consider RS regarding metric $v$ to be high when the model performs well in terms of $v$ (i.e., high mAP or low UQM) and the difference in $v$ between $X$ and $X$’s 10 adversarial examples is small. We then calculate $RS_{mAP}$ as:
$RS_{mAP}=Avg_{mAP}-Diff_{mAP},$ (14)
where $Avg_{mAP}$ is the mean value of mAP calculated by averaging the value achieved on $X$ and $X$’s 10 adversarial examples: $(mAP_{X}+\sum_{i=1}^{10}mAP_{X_{adv}^{i}})/11$; $Diff_{mAP}$ measures how differently the model performs on $X$ and $X$’s 10 adversarial examples in terms of mAP: $\sum_{i=1}^{10}|mAP_{X}-mAP_{X_{adv}^{i}}|/10$. As for $RS_{uq}$, we first calculate RS for each uncertainty metric $v$ as:
$RS_{v}=1-(Avg_{v}+Diff_{v}),$ (15)
where $v\in\{\textit{VR},\textit{SE},\textit{MI},\textit{TV},\textit{PS}\}$. $Avg_{v}$ and $Diff_{v}$ are calculated using the same equations as $Avg_{mAP}$ and $Diff_{mAP}$. We then calculate $RS_{uq}$ by taking the mean value of RS for each uncertainty metric:
$RS_{uq}=\frac{\sum_{v\in UQMs}RS_{v}}{count(UQMs)}.$ (16)
$RS_{uq}$ combines the RS of the five UQMs to show the overall robustness of the SDMs in terms of uncertainty.
Table 2: Datasets, metrics, and statistical tests for each RQ
RQ Dataset Metric Statistical Test
RQ1
origImg, dalleImg, stableImg
mAP Friedman test, Wilcoxon Signed-Rank test, rank-biserial correlation, Holm–Bonferroni method
RQ2
origImg, dalleImg, stableImg
VR, SE, MI,
TV, PS
RQ3
origImg, advOrigImg, dalleImg,
advDalleImg, stableImg, advStableImg
$RS_{mAP}$, $RS_{uq}$
RQ4
origImg, dalleImg, stableImg
mAP, VR, SE,
MI, TV, PS
Spearman’s rank correlation,
Holm–Bonferroni method
4.4 Research Questions
We answer the following four research questions (RQs): RQ1: How does each SDM perform when detecting stickers regarding detection accuracy? RQ2: How does each SDM perform when detecting stickers regarding prediction uncertainty? RQ3: How robust is each SDM in detecting stickers? RQ4: How does the detection accuracy correlate with the prediction uncertainty? Table 2 describes the employed datasets, metrics, and statistical tests for answering RQs.
4.5 Statistical Test
Since we compare more than two paired groups, i.e., SDMs using the same image dataset, we conduct the statistical analysis using the Friedman test [31], the Wilcoxon Signed-Rank test [32], and rank-biserial correlation [33]. Following the guideline [34, 35], we employ the Friedman test to verify if there are overall significant differences among all the paired groups (i.e., the models). If at least one group differs significantly from the others, $p<0.05$ will be computed by this statistical test. If this is the case, we apply the Wilcoxon Signed-Rank test for pairwise comparisons (i.e., comparing two models) with a significance level of 5%. To interpret the Wilcoxon Signed-Rank test more precisely, we calculate the median and the standard deviation by the median absolute deviation (MAD) of our data, i.e., [median $\pm$ 2 MAD] [36, 37]. It helps to check if the underlying distribution is stochastically equal to, less than, or greater than a distribution symmetric about zero, and accordingly define the relevant alternative hypothesis for more effective execution of the Wilcoxon Signed-Rank test. Moreover, for pairwise groups with significant differences, we use the rank-biserial correlation as the effect size of the Wilcoxon Signed-Rank test to determine which group shows better results. The rank-biserial correlation ranges from -1 to 1, where a positive result indicates that the first group tends to be larger than the second, and a negative result means the opposite, whereas 0 explains there is no significant difference between the two paired groups. Based on the guideline [38, 39], the effect size magnitude of the rank-biserial correlation involves four levels: negligible ($>$ 0.0 and $<$ 0.1), small ($\geq$ 0.1 and $<$ 0.3), medium ($\geq$ 0.3 and $<$ 0.5), and large ($\geq$ 0.5 and $\leq$ 1.0). We consider medium and large levels to demonstrate significant differences in our data.
To study the correlation between mAP and UQMs for RQ4, we use Spearman’s rank correlation coefficient [40], which provides the correlation coefficient ($\rho$) and the significance level ($p$). A $p$ less than 0.05 indicates a significant correlation between mAP and UQMs. Whereas correlation coefficient $\rho$ ranges from -1 to 1, revealing the direction and strength of the correlation. A positive $\rho$ indicates a positive correlation, meaning that one variable increases as the other increases, and vice versa. A $\rho$ of 0 shows no correlation.
Considering that the Wilcoxon Signed-Rank test and Spearman’s rank correlation coefficient are utilized multiple times to compare multiple paired groups, we address the issue of aggregated error probabilities resulting from multiple comparisons by applying the Holm–Bonferroni method [41]. This multiple comparison correction technique adjusts the significance level from the overall and controls the family-wise error rate at an $\alpha$ level, e.g., 5%.
5 Results and Analyses
We present descriptive statistics and summarize the key results of the statistical tests in Sections 5.1 to 5.4, and then provide guidelines for selecting SDMs in Section 5.5. The complete statistical results for each RQ are available in our replication package [42].
5.1 Results for RQ1 – Mean Average Precision
Figure 3 shows the mAP results achieved by each SDM for each dataset at different dropout rates with 95% confidence intervals. For the origImg dataset, Faster R-CNN, Faster R-CNN_v2, RetinaNet, and RetinaNet_v2 all achieved mAP values close to 1 across all dropout rates, which indicates these four SDMs achieve high precision and accuracy in detecting stickers correctly, and their performance is stable as the dropout rate increases. Lower mAP values are observed for SSD300 and SSDLite, with SSDLite consistently performing the worst. Moreover, the mAP values for both SDMs (SSD300 and SSDLite) decrease as the dropout rate increases, indicating that they are more susceptible to the dropout rate. For dalleImg dataset, Faster R-CNN_v2 achieved the consistent best performance at all dropout rates. Faster R-CNN, RetinaNet, and RetinaNet_v2 performed comparably, which slightly outperformed SSD300. SSDLite performed consistently the worst. For stableImg dataset, we can observe that at all dropout rates, Faster R-CNN_v2 performed the best followed by SSD300 and Faster R-CNN. SSDLite outperformed RetinaNet and RetinaNet_v2 at dropout rates below 0.2, after which the performance of SSDLite started to decrease significantly and underperformed RetinaNet and RetinaNet_v2, while RetinaNet and RetinaNet_v2 performed consistently across all dropout rates. Besides, when comparing the mAP achieved by SDMs on LLM-synthetic datasets (dalleImg and stableImg) and origImg, we observe that Faster R-CNN, Faster R-CNN_v2, RetinaNet, and RetinaNet_v2 perform poorer on the LLM-synthetic datasets than on origImg dataset, while SSD300 and SSDLite exhibit better performance on the LLM-synthetic datasets than on origImg dataset.
Figure 3: Average mAP achieved by SDMs for origImg, dalleImg, and stableImg at different dropout rates. M1: Faster R-CNN, M2: Faster R-CNN_v2, M3: RetinaNet, M4: RetinaNet_v2, M5: SSD300, and M6: SSDLiteRQ1.
We rank the best SDM that significantly outperformed the other SDMs on each metric according to the pairwise comparison results. For SDMs having no significant differences, we consider them to be a tie. Table 3 shows the results of best SDM(s) for each metric, where the last row shows the SDM recommendations based on different metrics. The best SDM for mAP is shown in the second column of Table 3. We can observe that for origImg dataset, Faster R-CNN, Faster R-CNN_v2, and RetinaNet_v2 all performed the best in terms of mAP, significantly outperforming the other SDMs (RetinaNet, SSD300, and SSDLite) across all the dropout rates. As for dalleImg and stableImg datasets, Faster R-CNN_v2 significantly outperforms the other five SDMs. When looking at the best SDM across all datasets, Faster R-CNN_v2 is recommended for all three datasets.
Conclusion for RQ1: Different SDMs perform differently in terms of mAP on different datasets. Specifically, Faster R-CNN, Faster R-CNN_v2, and RetinaNet_v2 all perform the best for origImg dataset, while Faster R-CNN_v2 achieves the best performance for dalleImg and stableImg datasets. Regarding the best SDM across all datasets, Faster R-CNN_v2 achieves the overall best performance regarding mAP and is recommended.
[…]
Dataset RQ1 RQ2 RQ3
mAP VR SE MI TV PS $\bm{\mathit{RS_{mAP}}}$ $\bm{\mathit{RS_{uq}}}$
origImg M1,2,4 M1,2,5,6 M1,4 M4 M4 M4 M1,4 M4
dalleImg M2 M1-6 M4 M4 M4 M4 M2 M4
stableImg M2 M1-6 M4 M4 M4 M4 M2 M4
$\bm{\mathit{Rec_{metric}}}$ M2 M1,2,5,6 M4 M4 M4 M4 M2 M4
5.2 Results for RQ2 – Uncertainty
Figure 4 presents the mean results regarding the five UQMs for each dataset at different dropout rates with 95% confidence intervals.
Uncertainty for Label Classification. We have the following observations regarding three label classification UQMs (i.e., VR, SE, and MI). For VR, Faster R-CNN, Faster R-CNN_v2, RetinaNet, RetinaNet_v2, SSD300, and SSDLite all achieve a very low VR, i.e., less than 0.015 for all three datasets at all dropout rates. The results suggest that all SDMs exhibit negligible uncertainty in label classification dispersion in the three datasets, that is, when an object is detected multiple times, the SDMs are confident that they will give the same classification result. However, for origImg dataset, we observe slightly higher VR values, which increase as the dropout rate increases for both RetinaNet and RetinaNet_v2, showing higher uncertainty and less stability of these two SDMs in terms of VR as dropout rate increases. As for SE and MI, SSDLite consistently achieves the highest uncertainty values of both UQMs for all datasets at all dropout rates, followed by SSD300, which exhibits high uncertainty values regarding these two UQMs for origImg and dalleImg datasets. However, exceptions can be observed for stableImg dataset, where SSD300 shows lower SE and MI values than Faster R-CNN and Faster R-CNN_v2 at low dropout rates and starts to show higher uncertainty as the dropout rate increases. Besides, Faster R-CNN, Faster R-CNN_v2, RetinaNet, and RetinaNet_v2 achieve comparable performance in terms of SE and MI for origImg dataset at all dropout rates, while when it comes to dalleImg and stableImg datasets, RetinaNet and RetinaNet_v2 consistently achieve the lowest SE and MI values. Recall that SE quantifies the amount of information in the predictions, while MI measures the difference in information between the prediction and the posterior. The results for SE and MI show that the prediction of SSD300 and SSDLite contain more information and differ more from the posterior, thus exhibiting higher uncertainty in label classification as compared to the other four SDMs, i.e., Faster R-CNN, Faster R-CNN_v2, RetinaNet, and RetinaNet_v2.
Figure 4: Average UQMs achieved by SDMs for origImg, dalleImg, and stableImg at different dropout rates. M1: Faster R-CNN, M2: Faster R-CNN_v2, M3: RetinaNet, M4: RetinaNet_v2, M5: SSD300, and M6: SSDLiteRQ2.
Uncertainty for Bounding Box Regression. Similar observations can be made for the two bounding box regressions UQMs (i.e., TV and PS), that is, SSD300 and SSDLite show higher TV and PS values for all datasets at all dropout rates except for TV on dalleImg dataset, where SSD300 achieves lower TV than Faster R-CNN until the dropout rate reaches 0.45. Besides, Faster R-CNN, Faster R-CNN_v2, RetinaNet, and RetinaNet_v2 achieve comparable and consistent low TV and PS, except for TV of Faster R-CNN on dalleImg dataset. The results of TV and PS show that SSD300 and SSDLite exhibit higher uncertainty in predicting bounding boxes than the other four SDMs. Moreover, when the dropout rate increases, the uncertainties in SSD300 and SSDLite also increase, while for the other SDMs, the dropout rate has little effect on their uncertainties.
Table 3 shows the best SDM(s), where we can find that regarding VR, Faster R-CNN, Faster R-CNN_v2, SSD300, and SSDLite achieve the best performance on origImg dataset, while all SDMs perform comparable on dalleImg and stableImg datasets. As for SE, MI, TV, and PS, RetinaNet_v2 perform the best, i.e., shows the lowest uncertainty when detecting stickers, except for SE on origImg dataset, where Faster R-CNN and RetinaNet_v2 achieve comparable best performance. When looking at the SDM recommendations across all datasets, regarding VR, Faster R-CNN, Faster R-CNN_v2, SSD300, and SSDLite are recommended, while for SE, MI, TV, and PS, RetinaNet_v2 is recommended as the best SDM.
Conclusion for RQ2: Different SDMs show different levels of uncertainties for different UQMs. Regarding VR, Faster R-CNN, Faster R-CNN_v2, SSD300, and SSDLite are recommended as the best SDMs since they have the lowest uncertainties across all datasets, while for SE, MI, TV, and PS, RetinaNet_v2 achieves the lowest uncertainties and is recommended as the best SDM.
Figure 5: Average $\bm{\mathit{RS_{mAP}}}$ and $\bm{\mathit{RS_{uq}}}$ achieved by SDMs for origImg, dalleImg, and stableImg at different dropout rates. M1: Faster R-CNN, M2: Faster R-CNN_v2, M3: RetinaNet, M4: RetinaNet_v2, M5: SSD300, and M6: SSDLiteRQ3.
5.3 Results for RQ3 – Robustness
Figure 5 shows the mean $RS_{mAP}$ and $RS_{uq}$ results with 95% confidence intervals. Regarding $RS_{mAP}$, we observe that Faster R-CNN_v2 performs the best for dalleImg and stableImg datasets at all dropout rates, while for origImg dataset, Faster R-CNN, Faster R-CNN_v2, RetinaNet, and RetinaNet_v2 all perform the best with a $RS_{mAP}$ close to 1. SSDLite consistently performs the worst for origImg and dalleImg datasets at all dropout rates, and performs worst on stableImg when the dropout rate is above 0.2. Besides, the increase in dropout rate has different effects on different SDMs, for example, the dropout rate increase leads to a decrease of $RS_{mAP}$ for SSDLite, while does not have much impact on Faster R-CNN, Faster R-CNN_v2, RetinaNet, and RetinaNet_v2. Regarding $RS_{uq}$, we observe that the results are consistent for all three datasets at all dropout rates, that is, RetinaNet_v2 achieves the best $RS_{uq}$, while SSDLite performs the worst. We also see a decrease in $RS_{uq}$ as the dropout rate increases. We also show the best SDM(s) in Table 3. For $RS_{mAP}$, Faster R-CNN and RetinaNet_v2 perform the best on origImg dataset, while Faster R-CNN_v2 perform the best on dalleImg and stableImg datasets. When looking at the best SDM across all datasets, Faster R-CNN_v2 is recommended. Regarding $RS_{uq}$, RetinaNet_v2 performs consistently the best across all three datasets and is therefore recommended.
Conclusion for RQ3: The six SDMs perform differently for different datasets in terms of $RS_{mAP}$ and $RS_{uq}$. Specifically, regarding $RS_{mAP}$, Faster R-CNN and RetinaNet achieve the best performance on origImg dataset, and Faster R-CNN_v2 perform the best on dalleImg and stableImg datasets. Regarding $RS_{uq}$, RetinaNet_v2 performs consistently the best across all three datasets.
5.4 Results for RQ4 – Correlation
Table 4 shows the results of Spearman’s rank correlation test for the correlation between mAP and the five UQMs for different SDMs. As for the correlation between mAP and VR, for all SDMs the difference is either insignificant or non-existent, while in terms of the correlation between mAP and the other four UQMs (SE, MI, TV, and PS), negative correlations can be observed for all SDMs, except for SSD300, which achieves insignificant correlation between mAP and TV/PS. This observation suggests that as uncertainty increases (except for VR), detection precision and accuracy will decrease. In addition, we also observe that as the dropout rate increases, the correlation between mAP and SE, MI, TV, and PS increases. This is reasonable since higher dropout rates will introduce higher uncertainties in the predictions.
Conclusion for RQ4: We observe significant negative correlations between mAP and SE, MI, TV, and PS for all SDMs, except for TV and PS achieved by SSD300, while insignificant or non-existent correlations are observed for the correlation between mAP and VR. Besides, the correlation between mAP and SE, MI, TV, and PS increases as the dropout rate increases.
Table 4: Results of the Spearman’s rank correlation coefficient between mAP and UQMs for each SDM on all dropout rates and datasets. A value or N/A denotes that the correlation is not statistically significant or non-existent, otherwise, it is significant, i.e., $\bm{p<0.05}$RQ4.
SDM VR SE MI TV PS
Faster R-CNN N/A -0.312 -0.277 -0.476 -0.461
Faster R-CNN_v2 N/A -0.408 -0.357 -0.365 -0.334
RetinaNet 0.012 -0.161 -0.107 -0.328 -0.329
RetinaNet_v2 0.098 -0.209 -0.065 -0.316 -0.295
SSD300 N/A -0.530 -0.455 -0.016 0.047
SSDLite N/A -0.510 -0.070 -0.269 -0.174
5.5 Concluding Remarks and Guidelines
We analyze the performance of each SDM in terms of detection accuracy (mAP), prediction uncertainty (UQM), and adversarial robustness ($RS_{mAP}$ and $RS_{uq}$). Based on the results, we observe that different SDMs perform differently regarding different metrics. In practice, considering different purposes and applications, selecting the appropriate SDM based on specific performance requirements is crucial. Therefore, we provide guidelines for selecting SDMs from various perspectives in Figure 6. Regarding detection accuracy, i.e., mAP, Faster R-CNN_v2 is recommended as it achieves the overall best performance. As for UQMs, for both label classification and bounding box regression tasks, we recommend RetinaNet_v2 because it shows lower uncertainty in terms of both types of uncertainty metrics. When it comes to the robustness of SDMs, Faster R-CNN_v2 shows the overall highest robustness in terms of detection accuracy, i.e., $RS_{mAP}$, and therefore recommended as the best SDM, while when considering robustness regarding uncertainty, RetinaNet_v2 is recommended as the best SDM.
Figure 6: Guidelines for Selecting Sticker Detection Models.
6 Threats to Validity
Conclusion Validity concerns the reliability of the conclusions. We employed appropriate statistical tests to draw reliable conclusions and followed a rigorous statistical procedure to analyze the collected data. Internal Validility is related to the parameter settings. To determine the hyperparameters for the clustering algorithm, i.e., HDBSCAN, we ran different combinations of hyperparameter values on a small subset of the original dataset and selected the combination that correctly clustered all objects. Besides, the threshold value $t$ for IoU is set to 0.5, following Catak et al. [20]. We, however, acknowledge that experimenting with various $t$ is needed to study its effect on the results, which is one of our further plans. Constructive Validity concerns the metrics used for the evaluation and we employed comparable metrics to ensure fair comparisons. Specifically, we used mAP to measure the sticker detection performance and five UQMs commonly used in classification and regression to measure model uncertainty. We also calculated RS to measure the robustness of the models under adversarial attacks. External Validility is about the generalizability of the empirical evaluation. We employ six models and six datasets. These models are based on different architectural designs and represent state-of-the-art object detection models. Regarding the benchmark datasets, in addition to the original dataset, we synthesize two datasets by prompting two state-of-the-art VLMs, i.e., DALL-E-3 and Stable Diffusion-3. Besides, we also generate adversarial datasets by perturbing the original and synthetic datasets.
7 Lessons Learned and Industrial Perspectives
Understanding, Utilizing, and Handling Uncertainty in Refurbishing Laptops. We quantify uncertainty in the SDMs from two aspects: uncertainty in label classification and bounding box regression, where the former tells the confidence of SDMs in the classification of detected stickers, whereas the latter represents their confidence and variability in localizing the positions of the detected stickers. UQ plays a pivotal role in enhancing the refurbishment of laptops by guiding laptop image data collection, ensuring dataset diversity, and determining the need for manual intervention. Collecting laptop images is time-consuming and significantly impacts solution quality, and UQ can guide decisions on whether additional data is needed. Specifically, UQ helps identify areas with high prediction uncertainty, indicating the need for more diverse and comprehensive training data. Besides, UQ is essential in addressing the limitations of existing datasets created by DTI, which primarily features Lenovo laptops and common sticker placements, e.g., stickers on the back cover. Also, images in the DTI dataset were captured in a relatively sterile setup with high quality, which may not reflect real-world conditions. Therefore, SDMs need to be tested under different conditions in terms of handling unexpected situations to ensure robustness, such as stickers on the side of a laptop or stickers near a mouse pad, where processor branding stickers are often found. By highlighting gaps in data diversity, UQ directs efforts towards collecting images from various laptop brands and unexpected sticker locations, ensuring SDMs are robust and adaptable to real-world conditions. Finally, UQ informs the necessity of manual intervention in the refurbishment process, enabling efficient and reliable operations by indicating when human oversight is required.
In addition to the uncertainty in SDMs, the overall process of laptop refurbishing experiences uncertainties from various aspects, such as humans working in collaboration with robots, software deployed in robots, hardware errors, and other environmental factors (e.g., lighting conditions). Hence, various UQ methods are needed to capture uncertainties in the whole process, not only related to the employed SDMs. As a result, an integrated UQ method is required for the entire laptop refurbishing process that quantifies uncertainties in various aspects and eventually provides an overall uncertainty associated with the whole process. Such quantified uncertainty can play a key role in identifying factors contributing to uncertainties, followed by devising best practices and methods to reduce the overall uncertainty.
Realism of VLM-Generated Images. To assess the SDMs, we generated images of laptops with stickers using two commonly used pre-trained VLMs. However, this poses an open question of whether the images generated are realistic. In our context, all the generated images were manually checked for realism. However, such an approach is not scalable, especially when generating a large-scale dataset. To this end, we foresee the need for an automated method to generate realistic images of laptops with stickers. Pre-trained VLMs already possess a strong understanding of general image features, and domain-specific fine-tuning can further refine their output. Therefore, one possible solution is to fine-tune VLMs to adapt to specific domains related to laptops and stickers, thereby helping produce more realistic images from the outset and reducing the need for extensive post-production filtering. Besides, large language models have been proven to be a way of realism evaluation [43]. Furthermore, training a realism classifier to distinguish realistic from unrealistic images can automate the manual check process. In the future, we plan to investigate scalable and efficient methods to generate large-scale realistic laptop image datasets.
8 Related Works
We discuss related works from two perspectives, i.e., UQ in DNNs and robustness assessment of DNNs.
UQ in DNNs. UQ helps enhance model reliability and trustworthiness. Various UQ methods have been studied to measure confidence in the model predictions. BNN provides a probabilistic framework for DNNs through Bayesian inference [5]. It is systematic but computationally expensive and has a complex implementation. As a Bayesian approximation, MC-Dropout is a practical and widely used UQ method [6], which reduces the computational cost. Deep Ensembles (DE) [44] involves independently training multiple DNNs with different initialization, and the variance in the predictions of the multiple DNNs serves as an uncertainty measure. DE is robust and reliable in uncertainty estimation but computationally and memory-wise expensive, requiring training and storing multiple models. UQ methods have been applied in many domains, e.g., cyber-physical systems [45, 46], computer vision tasks [47, 19, 48], and heathcare [49]. In contrast, this paper adopts MC-dropout as the UQ method to quantify the uncertainty of sticker detection tasks in laptop refurbishing robots, thereby studying their real-world application in a new context.
Robustness Assessment of DNNs. Robustness measures DNNs’ reliability. Recent robustness evaluation methods focus on adversarial robustness [7], which refers to the ability of DNNs to maintain performance and provide reliable outputs in the face of various perturbations, noises, and adversarial attacks. For instance, Carlini and Wagner [4] constructed three adversarial attack methods to evaluate the robustness of defensively distilled networks. Madry et al. [8] proposed a Projected Gradient Descent (PGD) attack to assess the adversarial robustness of DNNs from the robust optimization perspective. PGD employs testing accuracy as an estimation to evaluate the model’s robustness. Croce and Hein [50] further proposed two extensions of the PGD attack to address the problem of overestimating robustness in PGD. Yu et al. [51] qualitatively interpreted the adversarial vulnerability and robustness of DNNs and proposed a quantitative metric for the model-intrinsic robustness evaluation by measuring the model’s prediction divergence. Brendel et al. [52] combined decision-based boundary attack with the gradient-based adversarial attack to achieve fast and accurate robustness evaluation. This paper employs a dense adversarial attack technique to measure adversarial robustness. Regarding the robustness metrics, in addition to measuring the robustness using the performance metric (i.e., mAP), we further calculate the robustness score concerning UQMs.
9 Conclusions and Future Work
In this paper, we conduct an empirical study to evaluate the detection accuracy, prediction uncertainty, and adversarial robustness of six sticker detection models in the laptop refurbishment process. We adopt the Monte Carlo Dropout method to quantify the prediction uncertainty and measure uncertainty from two aspects: uncertainty in classification and regression. In addition, we present novel robustness metrics to evaluate the robustness of the models regarding detection accuracy and uncertainty. As for the benchmark datasets, in addition to the original dataset, we synthesize two datasets using two vision language models. Besides, we generate three adversarial datasets using an adversarial attack method. The results show different performances of different models, and Faster R-CNN_v2 and RetinaNet_v2 achieve the overall best performance in terms of all metrics. We also provide model selection guidelines and lessons learned. Future works include quantifying uncertainties of the overall laptop refurbishing process and investigating methods to generate large-scale realistic laptop image datasets.
Acknowledgments
The work presented here is supported by the RoboSapiens project funded by the European Commission’s Horizon Europe programme under grant agreement number 101133807.